# What makes unlearning hard and what to do about it

### v1

@@ 6 Experimenting with RUM flavours
We experiment with a refinement strategy based on memorization scores where $K$ = 3. Specifically, we study unlearning a forget set $\mathcal{S}$ that is the union of the three sets containing the $N$ lowest, the $N$ closest to 0.5, and the $N$ highest memorized examples in the dataset, where $N=1000$, making the overall size of the forget set 3000. We conduct experiments in two different datasets: CIFAR-10 and CIFAR-100, using ResNet-18 and ResNet-50, respectively. We evaluate unlearning algorithms both in terms of ToW, as before, and using a commonly-used MIA (Fan et al., 2023; Liu et al., 2024) that is a binary classifier trained to separate $\mathcal{R}$ from $\mathcal{D}_{test}$ and then queried on examples from $\mathcal{S}$. Following previous work, we report as “MIA” the fraction of examples from $\mathcal{S}$ that were predicted to be held-out. The goal is to match the “MIA” score of retrain-from-scratch. For convenience, we report the “MIA gap”: the absolute difference of the MIA score of unlearning from the MIA score of retrain-from-scratch (lower is better). See Section A.4 for details. We also consider and analyze the effect of two different orderings: i) low $\rightarrow$ medium $\rightarrow$ high and i) high $\rightarrow$ medium $\rightarrow$ low.
@@
To investigate this, we selected a subset of highest-performing algorithms and apply each algorithm $\mathcal{U}$ in three ways: i) “vanilla”, i.e. applying $\mathcal{U}(\theta^{o},\mathcal{S},\mathcal{R})$ as usual, ii) “shuffle” where $\mathcal{U}$ is applied sequentially on three equal-sized subsets of $\mathcal{S}$ that were determined randomly, and iii) RUMF, where $\mathcal{U}$ is applied sequentially on the subsets obtained by $\mathcal{F}(\mathcal{S})$, i.e. utilizing only the refinement step of RUM and applying the same $\mathcal{U}$ on each subset. We include “shuffle” as a control experiment, so that any gain of RUMF over shuffle is due to homogenization rather than simply reducing the size of the forget set or other effects of sequential unlearning. We observe from Figure 3 and Tables 11(a) and 11(b) that, for four different unlearning algorithms and on two different datasets, RUMF significantly outperforms “vanilla” and “shuffle”, indicating that operating sequentially on homogenized subsets according to memorization can boost the performance of unlearning algorithms. Interestingly, in most cases, “shuffle” actually performs worse than “vanilla”, indicating that the difficulty of the problem may increase rather than decrease given a poor refinement strategy.
@@
We report ToW and MIA with different orderings in Tables 11(a) and 11(b). We find that, while ToW is similar for different orderings, MIA can vary. To better understand these dynamics, we inspect the accuracies on $\mathcal{S}$ and its subsets after each step in Figure 4 for SalUnF on CIFAR-10 and in Figure 9 for NegGrad+F CIFAR-100. The former reveals why SalUnF with low $\rightarrow$ med $\rightarrow$ high order greatly outperforms vanilla SalUn. Recall that we identified in Section 4.2 that SalUn “overforgets” low-mem examples (its forget accuracy is lower than that of retraining). We observe from Figure 4 that future steps of the sequence neutralize that overforgetting effect on low-mem, leading to better ToW (see Figure 3). Interestingly, in line with our previous insights (Section 4.2), we find (from both Figures 4 and 9) that it is hard to cause the “low-mem” accuracy to become low and stay low. NegGrad+ does not drop it for any order of execution; SalUn drops it in the first ordering, but that drop is later reversed. We leave it to future work to further study these sequential dynamics and their helpful or harmful effects on Tow and MIA. The fact that MIA results differ based on the sequence may tie in with the recently-identified “privacy onion effect” Carlini et al. (2022b).
@@ 7 Discussion and conclusion
How is this important aspect affected in our sequential framework? We remark that it depends on the unlearning algorithm. For instance, applying Fine-tune three times is much more expensive than applying it once, because Fine-tune performs (at least) one epoch over the entire retain set. But for other algorithms the overall cost does not increase significantly. Further, a key observation from our results is that we can do well by actually doing nothing on a subset of the forget set, which can really boost efficiency (especially since the vast majority of examples are “low mem”).
@@
How does the embedding space entanglement interact with the level of memorization of the forget set? We analyzed this and found in Table 3 that all of our memorization buckets have relatively high ES, indicating that separating out (data-space) outliers in the forget set doesn’t lead to lower entanglement between the two sets in the embedding space. We leave it to future work to study the interaction of these factors.
@@
We hope future work explores other refinement strategies (e.g. for a notion that captures embedding-space entanglement), designing refinement and meta-unlearning strategies that are practical, e.g. leveraging cheap proxies for quantities like memorization scores, investigating privacy implications of sequential RUM, e.g. in terms of the “privacy onion effect” Carlini et al. (2022b). We also hope to see how our RUM framework can be adopted and adapted for unlearning in LLMs, especially given the findings from the contemporaneous paper Barbulescu and Triantafillou (2024) where unlearning is performed only on the highest-memorized examples in the forget set (albeit, memorization is defined differently for LLMs). We hope our framework continues to enable progress in understanding and improving unlearning and that our identified factors of difficulty and associated behaviours of existing algorithms continue to improve the state-of-the-art and inform the development of strong evaluation metrics that consider forget sets that vary in terms of relevant identified characteristics.
@@ A.2 Implementation details
We use CIFAR-10 and CIFAR-100 datasets for our evaluation. The CIFAR-10 dataset consists of 60,000 32x32 color images, including 10 classes of 6000 images each. The CIFAR-100 dataset has 100 classes, each containing 600 32x32 color images. In our experiments, the sizes of train, validation, and test sets are 45000, 5000, and 10000 respectively. We use ResNet-18 and ResNet-50 as model architectures. Specifically, ResNet-18 is used for CIFAR-10 and ResNet-50 is used for CIFAR-100, respectively.
@@
We trained a ResNet-18 model on the CIFAR-10 dataset and a ResNet-50 model on the CIFAR-100 dataset. For CIFAR-10, the ResNet-18 model was trained for 30 epochs using the SGD optimizer. The learning rate was initialized at 0.1 and scheduled with a cosine decay. We used a weight decay of 0.0005, a momentum of 0.9, and a batch size of 256. For CIFAR-100, the ResNet-50 model was trained for 150 epochs using the SGD optimizer. The learning rate was initialized at 0.1 and decayed by a factor of 0.2 at 60 and 120 epochs. We applied a weight decay of 0.0005, a momentum of 0.9, and a batch size of 256. Additionally, data augmentations, including random cropping and random horizontal flipping, were employed during training on CIFAR-100. All models were trained on Nvidia RTX A5000 GPUs.
@@ A.3 Procedure for creating retain / forget partitions with varying ES
This procedure enabled us to create retain / forget partitions with varying ES values. The ES values for our low, medium, and high ES partitions are shown in Table 2. As observed from the table, the ES values increase from low to high ES partitions for both CIFAR-10 and CIFAR-100, confirming the effectiveness of our procedure.
@@
Table 2: ES values for the low, medium and high forget / retain partitions for CIFAR-10 and CIFAR-100.
@@
A.4 Description of MIA
We adopted a MIA based on prediction confidence, following the procedure described by Liu et al. [2024]. To conduct this attack, we first sampled equal-sized data from the retain set and the test set, using these to train a binary classifier as the MIA model. This model is designed to distinguish whether a data example was involved in the training stage or not.
Next, we applied this attack model to the forget set to evaluate unlearning performance during the testing phase, after an unlearning method was implemented. Intuitively, for successful unlearning, we want forget set examples to be classified as “non-training” data. We define “training" data as the positive class and “non-training" data as the negative class, and measured the performance of MIA by calculating the ratio of true negatives (i.e., the number of the forgetting samples predicted as non-training examples) predicted by the MIA model to the total size of the forget set, as shown in (4), following the same procedure as prior work Liu et al. [2024].
@@ A.5 Data-space vs embedding-space outliers
We examined the ES values for each memorization category to understand the entanglement between the forget and retain sets in the embedding space. Table 3 presents the ES values corresponding to low, medium, and high memorization buckets. The findings reveal that all the memorization bucket exhibits relatively high ES, suggesting that separating outliers in the data space does not reduce the entanglement between the forget and retain sets in the embedding space.
@@
A.6 Detailed results

### v2

@@ 6 Experimenting with RUM flavours
We experiment with a refinement strategy based on memorization scores where $K$ = 3. Specifically, we study unlearning a forget set $\mathcal{S}$ that is the union of the three sets containing the $N$ lowest, the $N$ closest to 0.5, and the $N$ highest memorized examples in the dataset, where $N=1000$, making the overall size of the forget set 3000. We conduct experiments on two different datasets: CIFAR-10 and CIFAR-100, using ResNet-18 and ResNet-50, respectively. We evaluate unlearning algorithms both in terms of ToW, as before, and using a commonly-used MIA [10, 26] that is a binary classifier trained to separate $\mathcal{R}$ from $\mathcal{D}_{test}$ and then queried on examples from $\mathcal{S}$. Following previous work, we report as “MIA” the fraction of examples from $\mathcal{S}$ that were predicted to be held-out. The goal is to match the “MIA” score of retrain-from-scratch. For convenience, we report the “MIA gap”: the absolute difference of the MIA score of unlearning from the MIA score of retrain-from-scratch (lower is better). Further details on MIA setup are in Section A.4. We also introduce a new metric“ToW-MIA”, which mirrors ToW but replaces its "forget quality" component (i.e., its accuracy on $\mathcal{S}$) with the MIA score. Results for ToW-MIA are presented in Section A.4 and A.8 and largely follow the same trends. To gain a finer understanding of the differences between the unlearned and retrained models, we additionally analyze the number of examples where their predictions disagree (see Section A.7). We also consider and analyze the effect of two different orderings: i) low $\rightarrow$ medium $\rightarrow$ high and i) high $\rightarrow$ medium $\rightarrow$ low.
@@
To investigate this, we selected a subset of highest-performing algorithms and apply each algorithm $\mathcal{U}$ in three ways: i) “vanilla”, i.e. applying $\mathcal{U}(\theta^{o},\mathcal{S},\mathcal{R})$ as usual, ii) “shuffle” where $\mathcal{U}$ is applied sequentially on three equal-sized subsets of $\mathcal{S}$ that were determined randomly, and iii) RUMF, where $\mathcal{U}$ is applied sequentially on the subsets obtained by $\mathcal{F}(\mathcal{S})$, i.e. utilizing only the refinement step of RUM and applying the same $\mathcal{U}$ on each subset. We include “shuffle” as a control experiment, so that any gain of RUMF over shuffle is due to homogenization rather than simply reducing the size of the forget set or other effects of sequential unlearning. We observe from Figure 3 and Tables 16(a) and 16(b) that, for four different unlearning algorithms and on two different datasets, RUMF significantly outperforms “vanilla” and “shuffle”, indicating that operating sequentially on homogenized subsets according to memorization can boost the performance of unlearning algorithms. Interestingly, in most cases, “shuffle” actually performs worse than “vanilla”, indicating that the difficulty of the problem may increase rather than decrease given a poor refinement strategy.
@@
Can we obtain similar performance boosts with compute-efficient proxies?
Given the computational cost of calculating memorization scores, we aim to find a more efficient proxy to enable practical deployment of RUM. We discuss a confidence-based memorization metric, termed “C-proxy”, as a proxy for memorization [23, 32]. Section A.6 provides a detailed explanation of the proxy and the complete results across various datasets and architectures. Figure 4(a) shows that the observed unlearning difficulty pattern—specifically, that forget examples with higher memorization scores (i.e., lower C-proxy values, as they are negatively correlated; see Table 4) are harder to unlearn—remains consistent when using C-proxy, paralleling the results in Figure 1(b). This pattern is further confirmed on Tiny-ImageNet with both ResNet-18 and the VGG-16 architectures, as shown in Figure 9. We then examine whether using C-proxy achieves similar performance gains within RUM as the original memorization score. Figure 4(b) presents results on CIFAR-10 with ResNet-18, using C-proxy in place of memorization score during the refinement step. This analysis is also extended to Tiny-ImageNet with both ResNet-18 and VGG-16 architectures, as shown in Figure 10 and Table 17(c). Together, these results suggest that C-proxy is a practical and compute-efficient alternative, delivering significant performance gains comparable to those achieved with the memorization score. Detailed analysis of the use and appropriateness of various memorization proxies and their effect on RUM can be found in [37].
(a) C-proxy vs ToW
(b) RUMF on CIFAR-10 using C-proxy
Figure 4: Replacing mem scores with the efficient C-proxy yields similar trends and performance gains, carving a path for practical deployment of RUM. Left: forget sets with lower C-proxy values (i.e., higher mem scores, since C-proxy and memorization are negatively correlated) are harder to unlearn, consistent with the trend in Figure 1(b). Right: RUMF using C-proxy in the refinement step enhances unlearning performance across algorithms, comparable to using the memorization score in Figure 3(a). Error bars correspond to 95% confidence intervals, with each algorithm run 3 times.
@@
We report ToW and MIA with different orderings in Tables 16(a) and 16(b). We find that, while ToW is similar for different orderings, MIA can vary. To better understand these dynamics, we inspect the accuracies on $\mathcal{S}$ and its subsets after each step in Figure 5 for SalUnF on CIFAR-10 and in Figure 13 for NegGrad+F CIFAR-100. The former reveals why SalUnF with low $\rightarrow$ med $\rightarrow$ high order greatly outperforms vanilla SalUn. Recall that we identified in Section 4.2 that SalUn “overforgets” low-mem examples (its forget accuracy is lower than that of retraining). We observe from Figure 5 that future steps of the sequence neutralize that overforgetting effect on low-mem, leading to better ToW (see Figure 3). Interestingly, in line with our previous insights (Section 4.2), we find (from both Figures 5 and 13) that it is hard to cause the “low-mem” accuracy to become low and stay low. NegGrad+ does not drop it for any order of execution; SalUn drops it in the first ordering, but that drop is later reversed. We leave it to future work to further study these sequential dynamics and their helpful or harmful effects on Tow and MIA. The fact that MIA results differ based on the sequence may tie in with the recently-identified “privacy onion effect” [6].
@@ 7 Discussion and conclusion
How is this important aspect affected in our sequential framework? We remark that it depends on the unlearning algorithm. For instance, applying Fine-tune three times is much more expensive than applying it once, because Fine-tune performs (at least) one epoch over the entire retain set. But for other algorithms the overall cost does not increase significantly. Further, a key observation from our results is that we can do well by actually doing nothing on a subset of the forget set, which can really boost efficiency (especially since the vast majority of examples are “low mem”). Additionally, our results with the C-proxy demonstrate that significant performance improvements in unlearning can be achieved with minimal computational cost, avoiding the heavy expense of computing memorization scores.
@@
How does the embedding space entanglement interact with the level of memorization of the forget set? We analyzed this and found in Table 3(b) that all of our memorization buckets have relatively high ES, indicating that separating out (data-space) outliers in the forget set doesn’t lead to lower entanglement between the two sets in the embedding space. We leave it to future work to study the interaction of these factors.
@@
We hope future work explores other refinement strategies (e.g. for a notion that captures embedding-space entanglement) and investigates privacy implications of sequential RUM, e.g. in terms of the “privacy onion effect” [6]. We also hope to see how our RUM framework can be adopted and adapted for unlearning in LLMs, especially given the findings from the contemporaneous paper [2] where unlearning is performed only on the highest-memorized examples in the forget set (albeit, memorization is defined differently for LLMs). We hope our framework continues to enable progress in understanding and improving unlearning and that our identified factors of difficulty and associated behaviours of existing algorithms continue to improve the state-of-the-art and inform the development of strong evaluation metrics that consider forget sets that vary in terms of relevant identified characteristics.
@@ A.2 Implementation details
We use the CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets for our evaluation. CIFAR-10 consists of 60,000 32x32 color images across 10 classes, with 6,000 images per class. CIFAR-100 contains 100 classes, each with 600 32x32 color images. Tiny-ImageNet includes 100,000 64x64 color images across 200 classes, with each class containing 500 training images, 50 validation images, and 50 test images (for a total of 100,000 training, 10,000 validation, and 10,000 test images). For CIFAR-10 and CIFAR-100, the train, validation, and test set sizes are 45,000, 5,000, and 10,000 images, respectively. We use ResNet-18, ResNet-50, and VGG-16 as model architectures. Specifically, we use ResNet-18 for CIFAR-10, ResNet-50 for CIFAR-100, and both ResNet-18 and VGG-16 for Tiny-ImageNet.
@@
We trained four models across different datasets and architectures: ResNet-18 on CIFAR-10, ResNet-50 on CIFAR-100, ResNet-18 on Tiny-ImageNet, and VGG-16 on Tiny-ImageNet. For CIFAR-10, the ResNet-18 model was trained for 30 epochs using the SGD optimizer. The learning rate was initialized at 0.1 and scheduled with cosine decay. For CIFAR-100, the ResNet-50 model was trained for 150 epochs using the SGD optimizer, with the learning rate initialized at 0.1 and decayed by a factor of 0.2 at epochs 60 and 120. For Tiny-ImageNet, the ResNet-18 model was trained for 80 epochs, and the VGG-16 model was trained for 100 epochs, both using the SGD optimizer with an initial learning rate of 0.1 and cosine decay. All models were trained with a weight decay of 0.0005, a momentum of 0.9, and a batch size of 256. Additionally, data augmentations, including random cropping and random horizontal flipping, were applied during training on CIFAR-100 and Tiny-ImageNet. All training was conducted on Nvidia RTX A5000 GPUs.
@@ A.3 Procedure for creating retain / forget partitions with varying ES
This procedure enabled us to create retain / forget partitions with varying ES values. The ES values for our low, medium, and high ES partitions are shown in Table 2(b). As observed from the table, the ES values increase from low to high ES partitions for both CIFAR-10 and CIFAR-100, confirming the effectiveness of our procedure. We also use Maximum Mean Discrepancy (MMD) [18], a widely-used metric, to further validate ES. The MMD is computed as follows: Given a pre-trained model $\theta^{o}$ (the original model in our case) trained on $\mathcal{D}_{train}$, we first extract features for the forget set $\mathcal{S}$ and the retain set $\mathcal{R}$ using $\theta^{o}$, denoted as $\phi(\mathcal{S})$ and $\phi(\mathcal{R})$, respectively. We then apply an RBF kernel to map these features to another space $\mathcal{H}$ such that $\phi^{\prime}(\mathcal{S}),\phi^{\prime}(\mathcal{R})\in\mathcal{H}$. The MMD between $\phi(\mathcal{S})$ and $\phi(\mathcal{R})$ is then calculated using the following formula:
$\text{MMD}^{2}(\phi(\mathcal{S}),\phi(\mathcal{R}))=\left\|\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\phi^{\prime}_{i}-\frac{1}{|\mathcal{R}|}\sum_{j\in\mathcal{R}}\phi^{\prime}_{j}\right\|_{\mathcal{H}}^{2}$
Our experiments reveal a negative correlation between ES and MMD scores, as reported in Table 2(b). For CIFAR-10/ResNet-18 and CIFAR-100/ResNet-50, low (high) ES values correspond to high (low) MMD values, supporting the consistency of the ES.
@@
Table 2: ES and MMD scores for the low, medium and high forget / retain partitions for CIFAR-10 and CIFAR-100.
@@
A.4 Description of MIA and ToW-MIA
We adopted a MIA based on prediction confidence, following the procedure described by [26]. To conduct this attack, we first sampled equal-sized data from the retain set and the test set, using these to train a binary classifier as the MIA model. This model is designed to distinguish whether a data example was involved in the training stage or not.
Next, we applied this attack model to the forget set to evaluate unlearning performance during the testing phase, after an unlearning method was implemented. Intuitively, for successful unlearning, we want forget set examples to be classified as “non-training” data. We define “training" data as the positive class and “non-training" data as the negative class, and measured the performance of MIA by calculating the ratio of true negatives (i.e., the number of the forgetting samples predicted as non-training examples) predicted by the MIA model to the total size of the forget set, as shown in (3), following the same procedure as prior work [26].
@@ A.4 Description of MIA and ToW-MIA
Building on the MIA setup above, we introduce another metric “ToW-MIA” to evaluate unlearning performance, similar to ToW described in Section 4. The metric is defined as follows:
[…]
where $\text{m}(\theta,\mathcal{D})$ represents the MIA performance of the model $\theta$ trained on $\mathcal{D}$, as defined in Equation (3). The first term, $\text{dm}(\theta_{u},\theta_{r},\mathcal{S})=|\text{m}(\theta_{u},\mathcal{S})-\text{m}(\theta_{r},\mathcal{S})|$, denotes the absolute difference in MIA performance on $\mathcal{S}$ between the unlearned model $\theta_{u}$ and the retrained-from-scratch model $\theta_{r}$. This term is the key distinction between ToW and ToW-MIA, where “forget quality” is measured by MIA (ToW-MIA) rather than accuracy (ToW). The other two terms remain identical to those in ToW (see the equation in Section 4), representing the absolute difference in accuracy between $\theta^{u}$ and $\theta^{r}$ on $\mathcal{R}$ and $\mathcal{D}_{test}$, respectively, which reflect “model utility”. Like ToW, ToW-MIA rewards unlearned models that approximate the performance of the retrained model across the forget, retain, and test sets. ToW-MIA ranges from 0 to 1, with higher values indicating better unlearning performance. Figure 7 show the same trends with ToW-MIA as for ToW, both for unlearning difficulty analysis and RUM. More detailed results using ToW-MIA can be found in Table 14(c).
(a) ToW-MIA vs ES
(b) ToW-MIA vs memorization
(c) ToW-MIA vs RUMF
Figure 7: Evaluating performance with ToW-MIA (higher is better) for unlearning difficulty analysis (a, b) and RUMF (c). Error bars represent 95% confidence intervals from 3 runs per algorithm.
@@ A.5 Data-space vs embedding-space outliers
We examined the ES values for each memorization category to understand the entanglement between the forget and retain sets in the embedding space. Table 3(b) presents the ES values corresponding to low, medium, and high memorization buckets. The findings reveal that all the memorization bucket exhibits relatively high ES, suggesting that separating outliers in the data space does not reduce the entanglement between the forget and retain sets in the embedding space.
@@
A.6 Confidence-based proxy for memorization
To improve the efficiency of calculating memorization scores, we leverage insights from previous works [23, 32] and introduce a model confidence-based metric, referred to as the C-proxy, which serves as a proxy for the memorization score. The computation of the C-proxy proceeds as follows: For each data point $(x_{i},y_{i})\in\mathcal{D}$, we track the softmax probability of the model’s prediction $\theta(x_{i})$ for the ground-truth label $y_{i}$ across all training epochs, as the model $\theta$ is trained on $\mathcal{D}$ using algorithm $\mathcal{A}$. These probabilities are then averaged over all epochs at the end of training to capture the training dynamics.
Table 4 presents the fidelity (measured by Spearman correlation) and efficiency (measured by additional computational cost) of the C-proxy in relation to memorization. This evaluation is conducted on both CIFAR-10 and CIFAR-100 datasets. As shown, the C-proxy demonstrates a strong negative Spearman correlation with memorization scores while significantly reducing computational overhead compared to both computing memorization scores and retraining the model from scratch. This suggests that the C-proxy is an effective proxy for memorization while also offering substantial efficiency gains.
Table 4: Comparison of C-proxy based on Spearman correlation with memorization and relative computation time percentages in comparison to memorization computation / retraining the model from scratch, evaluated on CIFAR-10 and CIFAR-100 using ResNet-18 and ResNet-50, respectively.
Dataset Spearman corr. (mem) Comp. time % (mem) Comp. time % (retrain)
CIFAR-10 -0.80 0.018% 17.123%
CIFAR-100 -0.91 0.002% 8.175%
Figure 4(a) and Figure 9 present results across different datasets and model architectures, demonstrating that the pattern identified in Section 4.2 remains consistent when using the C-proxy instead of the memorization score. Specifically, the lower the C-proxy value, the harder it is to unlearn (note that the C-proxy is negatively correlated with memorization). Furthermore, Figure 4(b) and Figure 10 show the results of RUMF across various datasets and model architectures, using the C-proxy in place of memorization for the refinement step described in Section 5. More detailed results, including accuracy and MIA, are available in Table 17(c). These results demonstrate that RUMF still achieves a significant unlearning performance gain when C-proxy is used, comparable to those obtained with the memorization score. A detailed analysis of various memorization proxies, their suitability, and their impact on RUM can be found in [37].
(a) Tiny-ImageNet with ResNet-18
(b) Tiny-ImageNet with VGG-16
Figure 9: Impact of C-proxy on unlearning difficulty: C-proxy vs ToW on Tiny-ImageNet across low, medium, and high C-proxy levels, using ResNet-18 and VGG-16. Error bars represent 95% confidence intervals from 3 runs per algorithm.
(a) Tiny-ImageNet with ResNet-18
(b) Tiny-ImageNet with VGG-16
Figure 10: Impact of C-proxy on unlearning performance on Tiny-ImageNet with ResNet-18 and VGG-16. Each unlearning algorithm is applied in three ways: RUMF, as well as vanilla and shuffle as comparison. Error bars represent 95% confidence intervals from 3 runs per algorithm.
A.7 Per-example differences between the unlearned and retrained models
While accuracy is a commonly used metric in the unlearning literature [29], it reflects an average over a data distribution and may overlook finer details. To gain a more granular understanding of the similarities/differences between the unlearned model and the retrained model, we examine their behavior at the level of individual examples. Specifically, we collect predictions from both models for each example and count the examples where their predictions differ. Table 5(b) presents these results for a representative setting: CIFAR-10 dataset with ResNet-18 architecture.
We observe that, in all cases, the percentage of disagreements between the unlearned and retrained models remains low. For highly memorized data (which typically makes unlearning harder), the disagreement rate increases but stays modest, ranging between ca. [7%, 15%]. For low-memorized data, the disagreement rate is lower, within ca. [3%, 5%].
Interestingly, for the vast majority of unlearning algorithms, these per-example disagreement trends align with those observed using ToW: low-memorized examples are generally easier to unlearn, as indicated by both ToW and the lower disagreement rate, while high-memorized examples are harder to unlearn, with higher disagreement rates. A similar trend is observed across different ES levels as well. This consistency between average accuracies (ToW) and per-example disagreements supports the robustness of our findings and underscores the validity of ToW as a measure of unlearning difficulty.
Table 5: ToW vs percentage of different predictions across ES or memorization levels on CIFAR-10/ResNet-18, averaged over 3 runs with 95% confidence intervals.
ToW Percentage of different predictions (%)
Low ES Medium ES High ES Low ES Medium ES High ES
Original 0.944 $\pm$ 0.014 0.928 $\pm$ 0.022 0.759 $\pm$ 0.048 0.329 $\pm$ 0.069 0.396 $\pm$ 0.048 1.514 $\pm$ 0.212
Fine-tune 0.952 $\pm$ 0.024 0.923 $\pm$ 0.019 0.908 $\pm$ 0.018 0.624 $\pm$ 1.339 1.676 $\pm$ 2.431 3.218 $\pm$ 1.597
L1-sparse 0.945 $\pm$ 0.013 0.926 $\pm$ 0.019 0.836 $\pm$ 0.031 0.335 $\pm$ 0.051 0.407 $\pm$ 0.053 6.367 $\pm$ 11.164
NegGrad 0.929 $\pm$ 0.030 0.887 $\pm$ 0.047 0.766 $\pm$ 0.042 0.479 $\pm$ 0.413 3.853 $\pm$ 4.276 3.594 $\pm$ 3.802
NegGrad+ 0.941 $\pm$ 0.029 0.920 $\pm$ 0.038 0.800 $\pm$ 0.029 0.370 $\pm$ 0.022 1.796 $\pm$ 1.510 1.752 $\pm$ 0.174
Influence unlearning 0.928 $\pm$ 0.023 0.890 $\pm$ 0.012 0.744 $\pm$ 0.029 0.362 $\pm$ 0.208 0.986 $\pm$ 1.223 3.381 $\pm$ 3.342
SalUn 0.783 $\pm$ 0.031 0.656 $\pm$ 0.046 0.618 $\pm$ 0.056 8.816 $\pm$ 1.756 1.752 $\pm$ 5.706 5.706 $\pm$ 4.279
Random-label 0.767 $\pm$ 0.107 0.663 $\pm$ 0.055 0.443 $\pm$ 0.087 5.566 $\pm$ 0.406 12.154 $\pm$ 6.952 0.579 $\pm$ 0.579
(a) Per-example difference vs ES
ToW Percentage of different predictions (%)
Low mem Medium mem High mem Low mem Medium mem High mem
Original 0.988 $\pm$ 0.007 0.723 $\pm$ 0.053 0.231 $\pm$ 0.058 3.611 $\pm$ 0.856 5.081 $\pm$ 0.550 7.341 $\pm$ 0.822
Fine-tune 0.933 $\pm$ 0.052 0.884 $\pm$ 0.019 0.760 $\pm$ 0.065 5.962 $\pm$ 1.525 6.359 $\pm$ 1.135 7.099 $\pm$ 2.370
L1-sparse 0.914 $\pm$ 0.061 0.816 $\pm$ 0.011 0.629 $\pm$ 0.087 6.668 $\pm$ 1.676 10.163 $\pm$ 3.705 9.834 $\pm$ 2.030
NegGrad 0.938 $\pm$ 0.028 0.738 $\pm$ 0.005 0.325 $\pm$ 0.098 5.412 $\pm$ 1.020 9.073 $\pm$ 1.283 14.256 $\pm$ 3.437
NegGrad+ 0.965 $\pm$ 0.020 0.831 $\pm$ 0.032 0.661 $\pm$ 0.082 4.288 $\pm$ 0.452 6.357 $\pm$ 1.023 12.190 $\pm$ 2.720
Influence unlearning 0.986 $\pm$ 0.031 0.738 $\pm$ 0.051 0.381 $\pm$ 0.037 3.964 $\pm$ 0.803 9.668 $\pm$ 2.469 18.484 $\pm$ 1.491
SalUn 0.774 $\pm$ 0.043 0.758 $\pm$ 0.053 0.886 $\pm$ 0.082 5.963 $\pm$ 1.281 7.507 $\pm$ 0.268 8.031 $\pm$ 0.540
Random-label 0.709 $\pm$ 0.029 0.548 $\pm$ 0.093 0.877 $\pm$ 0.064 13.491 $\pm$ 5.363 12.973 $\pm$ 1.635 12.905 $\pm$ 4.413
(b) Per-example difference vs memorization
A.8 Detailed results